OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (7, –4); m = 5/3

OpenStudy (anonymous):

(y+4)/(x-7) = 5/3 3y+12= 5x-35 5x-3y-47=0